Logbooks general detector daq Analysis Physics 12 GeV HallA OldDVCS2 DVCS2Ana Parity
runliste issues procurement survey minutes web compton2 NewIssues NewProcurement
  Minutes  Not logged in ELOG logo
Message ID: 4     Entry time: 06/11/04, 18:06
Author: Franck Sabatie 
Subject: DVCS meeting minutes - 06/09/2004 
  Present: P. Bertin, C.E. Hyde-Wright, G. Gavalian, K. McCormick, M.
Mazouz, D. Hayes, A. Camsonne, C. Ferdi, A. Chantelauze, F. Sabatie

  Franck: Discussed the installation meeting with everybody in more
details. The installation schedule is mostly constrained by the first two
weeks while we wait for septa to be desinstalled. There will only be 7-10
days to check-out the detectors in the final configuration before Hall
locked period (starting Sept. 10th). This Hall locked period is still
unclear. We will probably be able to access the Hall in controlled access
if needed.

  David: showed the PA/tagger cabling scheme (trigger and ARS). Charles
and David payed attention to not get PA blocks and corresponding tagger
paddles in the same ARS module because of cross-talk. Attention was not
paid particularly to block-transfer issues since 0-suppression mode will
be chosen and will probably sparcify the data. David used a reversed
numbering scheme as we used in the Monte-Carlo. We decided to use MC
scheme so he will update his slides and make them available.

  Astrid: Details about LED system check-out using charles' light guide.
It turns out the system is very dependent on the alignment and very little
photons make it to the block. More details are available in Astrid's
recent ELOG entry.

  Alex: worked with M. Brossard on the new logic unit. Now operationnal.
Many scalers can be programmed in this unit (it contains a FPGA). Still a
few things to tune in DAQ. New trigger operationnal.

  Gagik: Worked on software a bit. Found a way to simplify library
structure. Will check in changes in CVS.
ELOG V3.1.4-unknown